The nucleus of an atom can be assumed to spheical.the radius of the nucleus of mass number a is given by 1.25 into 10to the power -13cm. radius of atom is one angstom. if the mass number is 64,then the fraction of the atomic volume that is occupied by the nucleus is
A)1.0 INTO 10 TO THE POWER-3
B)5.0 INTO 10 TO THE POWER-5
C)1.25 INTO 10 TO THE POWER -13

Grade:12

1 Answers

Pratham Ashish
14 years ago

hi ,

 radius of nucleus , R = (1.25 X10 ^-13 )* A ^1/3

                              =     (1.25 X10 ^-13 )* 64 ^1/3

   volume of nucleus  V =  4/3 ¶ { (1.25 X10 ^-13 )* 64 ^1/3 } ^3

                                  = 4/3 ¶ * 1.25 X 10 ^ -37   c.m.^3
       
    volume of atom =    4/3 ¶   ( 1.0 X 10 ^-8) ^3
             
                                = 4/3   ¶    x 10 ^ -24  c.m. ^3
 
   desired fraction =  4/3 ¶ * 1.25 X 10 ^ -37    / 4/3   ¶    x 10 ^ -24  
                              =  1.25 X 10 ^-13      C )